# Dynamics of interacting bosons in a two-leg ring ladder with artificial magnetic flux and ac-driven modulations

### v1

@@ IV Summary and Outlook
The application of an ac-driving induces particle excitations, where the distinct spectra are determined by the interplay between drive frequency, interaction strength, and local energy gap. When a finite inter-ring coupling is present and the intra-ring hopping is biased, the system exhibits chiral (currents flow in opposite directions) and antichiral dynamics (currents flow in the same direction), and the intensity and direction of particle currents can be precisely determined by jointly modulating the drive frequency, Peierls phase and interaction strength, which stimulates the chiral-antichiral transition.
This specific geometry can be experimentally realized with ultracold bosonic atoms in ring-shaped optical lattices, with the tunable tunneling and on-site energies being implemented via spatial optical modulators and artificial gauge fields [44, 45]. One could naturally extend the system to more complex configurations, such as multi-leg ring networks and two-dimensional lattices, and explore richer dynamical phases including topological pumping and quantum Hall-like behaviors. Moreover, the precise control of persistent currents and the transition between chiral and antichiral transports may contribute to quantum information transfer, novel atomtronic circuits, and topological quantum simulations.

### v2

@@ IV Summary and Outlook
The application of an ac driving induces particle excitations, where the distinct spectra are determined by the interplay between drive frequency, interaction strength, and local energy gap. When a finite inter-ring coupling is present and the intra-ring hopping is biased, the system exhibits chiral (currents flow in opposite directions) and antichiral dynamics (currents flow in the same direction), and the intensity and direction of particle currents can be precisely determined by jointly modulating the drive frequency, Peierls phase and interaction strength, which stimulates the chiral-antichiral transition.
This specific geometry can be experimentally realized with ultracold bosonic atoms in ring-shaped optical lattices, with the tunable tunneling and on-site energies being implemented via spatial optical modulators and artificial gauge fields [47, 62]. A weak local trap may be applied to the central sites of the rings before the atoms are loaded, such that they would be initially localized, and one subsequently switches off the trap, releasing the particles to evolve under the ac driving and modulating the interactions via Feshbach resonance. One could naturally extend the system to more complex configurations, such as multileg ring networks and two-dimensional lattices, and explore richer dynamics including topological pumping, quantum Hall-like behaviors and superfluid qubit systems [44, 65]. Moreover, the precise control of persistent currents and the transition between chiral and antichiral transports may contribute to quantum information transfer [44, 46], novel atomtronic circuits [66] and topological quantum simulations [67, 68].